OpenStudy (anonymous):

what is the taylor series representation for ln(1-x)?

OpenStudy (anonymous):

\[\ln (1-x)\]

OpenStudy (ghazi):

\[-x-\frac{ x^2 }{ 2 }-\frac{ x^3 }{ 3 }-\frac{ x^4 }{ 4 }-\frac{ x^5 }{ 5 }-.......\]

OpenStudy (anonymous):

how would you write it like a series? like with the sum symbol

OpenStudy (ghazi):

\[f(a)+\frac{ f^1(a) }{ 1! }(x-a)+\frac{ f^2(a) }{ 2! }(x-a)^2+\frac{ f^3(a) }{ 3! }(x-a)^3......\]

OpenStudy (ghazi):

use this formula and i just solved that and mentioned the answer